Grade 12th passMechanics

under constant pressure head the rate of flow of liquid through a cappiliary tube is V if length of the capillary is doubled & diameter of the bore is halved the rate of flow would become

akash
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Akash
 
From Poisson's formula
Rate of flow (V) is proportional to r⁴/l
V1/V2 = (r⁴/l)/ (r⁴/32l)
V2 = V1/32 = V/32
 
Regards
Arun (askIITians forum expert)